# Black holes in scalar multipolar Universes

### v1

@@ I Introduction
The Melvin magnetic Universe is an everywhere smooth, non-asymptotically flat, solution to the Einstein-Maxwell theory [2, 3]. It lends itself to many applications and generalizations ($e.g.$ [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]), including accommodating a Schwarzschild [15] (or Reissner-Nordström [16] or Kerr [17]) black hole (BH).
@@ IV.1 $a_{0}$: (decaying) monopole
This solution represents a Schwarzschild mass inside a scalar version of the Chazy-Curzon particle. Observe that even though in the auxiliary space we have chosen a single monopole, in spacetime there is a non-trivial multipolar dependence. Moreover, the would be horizon is now singular, as can be seen from the Ricci scalar
@@ IV.2 $b_{1}$: (growing) dipole
$\left(\frac{dr}{d\lambda}\right)^{2}=e^{b_{1}^{2}f(r)r^{2}\sin^{2}\theta}\left[E^{2}-\frac{j^{2}}{r^{2}}f(r)\right]\ .$ (28)
@@ V The Kaluza-Klein oxidization
$\mathcal{S}=\frac{1}{16\pi G_{5}}\int d^{4}\hat{X}\sqrt{-\hat{g}}\hat{R}\ ,$ (39)
@@ VI Discussion and further generalizations
$\mathcal{S}=\frac{1}{16\pi}\int d^{4}x\sqrt{-g}\left(R-\frac{1}{2}F_{\mu\nu}F^{\mu\nu}\right)\ ,$ (47)

### v2

@@ I Introduction
The Melvin magnetic Universe is an everywhere smooth, non-asymptotically flat, solution to the Einstein-Maxwell theory [2, 3]. It lends itself to many applications and generalizations ($e.g.$ [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]), including accommodating a Schwarzschild [16] (or Reissner-Nordström [17] or Kerr [18]) black hole (BH).
@@ IV.1 $a_{0}$: (decaying) monopole
This solution represents a Schwarzschild mass inside a scalar version of the Chazy-Curzon particle. Observe that even though in the auxiliary space we have chosen a single monopole, in spacetime there is a non-trivial multipolar dependence (for the scalar field). Moreover, the would be horizon is now singular, as can be seen from the Ricci scalar
@@ IV.2 $b_{1}$: (growing) dipole
$\left(\frac{dr}{d\lambda}\right)^{2}=e^{b_{1}^{2}f(r)r^{2}}\left[E^{2}-\frac{j^{2}}{r^{2}}f(r)\right]\ .$ (28)
@@ V The Kaluza-Klein oxidization
$\mathcal{S}=\frac{1}{16\pi G_{5}}\int d^{5}\hat{X}\sqrt{-\hat{g}}\hat{R}\ ,$ (39)
@@ VI Discussion and further generalizations
$\mathcal{S}=\frac{1}{16\pi}\int d^{4}x\sqrt{-g}\left(R-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}\right)\ ,$ (47)